Given: x = e2t - 1 and y = e2t + 1. Find y’.


1
Expert's answer
2022-02-27T12:20:03-0500

"x=e^{2t}-1"


"=> \frac{dx}{dt}=2e^{2t}"


"y=e^{2t}+1"


"=> \frac{dy}{dt}=2e^{2t}"


"y'=\frac{dy}{dx}=\frac{dy}{dt}×\frac{dt}{dx}"


"=> y'=\frac{dy}{dx}=\frac{2e^{2t}}{1}×\frac{1}{2e^{2t}}"



"=> y'=\frac{dy}{dx}=1"